OpenStudy (anonymous):

the equations of four lines are given. identify which lines are parallel Line 1: y=-7x+6 Line 2: x+1/5y=-6 Line 3: y=-5x-8 Line 4: y+7=-1/7(x+4)

OpenStudy (anonymous):

@pooja195 @Vocaloid

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@Pandabear78963 To know whether two lines are parallel is to see if they have the same slope, So its best to rewrite equations into slope-intercept form which is \(y=mx + b\) where \(m~ is~ slope\) and b is \(y~ intercept \) i tried like this... x +1/5y=-6 1/5y=-x - 6 then divide both sides by 1/5 y==-5x - 6 it too has a slope of -5 so they're parallel Did you get same answer?